OpenStudy (anonymous):

Sketch the region enclosed by the curves given below. Decide whether to integrate with respect to x or y. Then find the area of the region. 2 y = 5sqrt{x} , y = 3, 2 y + 2 x = 7 Area =

OpenStudy (anonymous):

How are you supposed to find the intersections?

OpenStudy (tkhunny):

These? \(2y = 5\sqrt{x}\) y = 3 2y + 2x = 7

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Find intersections by taking two at a time and solving. \(2y=5\sqrt{x}\) \( y = 3\) We have y = 3. We just need x. 2(3) = 5sqr(x) ==> sqrt(x) = 6/5 ==> x = 36/25 You do the next pair. \(2y=5\sqrt{x}\) \(2y + 2x = 7\)

OpenStudy (anonymous):

thats the one i am having trouble for

OpenStudy (anonymous):

\[2 y=5 \sqrt{x},4y^2=25x,y^2=\frac{ 25 }{ 4 }x\] it is a parabola. \[x=\frac{ 4 }{ 25 }y^2,2y+2*\frac{ 4 }{25 }y^2=7,8y^2+50y=175\] \[8y^2+50y-175=0,y=\frac{ -50 \pm \sqrt{(50)^2-4*8*-175} }{2*8 }\] \[y=\frac{ -50\pm \sqrt{2500+5600} }{ 16 }=\frac{ -50\pm \sqrt{8100} }{ 16 }=\frac{ -50\pm 90 }{ 16 }=\frac{ 40 }{ 16 },\frac{ -140 }{16 }\] \[y=\frac{ 5 }{2 },\frac{ -35 }{4 }\] \[when y=\frac{ 5 }{ 2 }, as given y is positive  \[required~ area=\int\limits_{5/2}^{3} x~dy\] \[=\int\limits_{5/2}^{3}\frac{ 7-2y }{ 2 }dy\] calculate.

OpenStudy (anonymous):

oh wow thnks

OpenStudy (anonymous):

but i am still troubled with how you got the intersections part.....